OpenStudy (anonymous):

2x^3y^4z^3a^2 over 4x^2y^5za Answer choices: A) axz over 2y B) xz^2a over 2y C) ax^2z over 2y D) xy over 2

OpenStudy (kinggeorge):

Just give it a shot. What do you think \[\large \frac{y^4}{y^5}\]simplifies as?

OpenStudy (anonymous):

y1

OpenStudy (kinggeorge):

Close, you have the right idea though. It should be \[y^{-1}=\frac{1}{y}\]

OpenStudy (anonymous):

o okay i was close

OpenStudy (kinggeorge):

Let's do another one. Can you simplify \[\large \frac{z^3}{z}?\]Just keep using the same idea.

OpenStudy (anonymous):

1^3?

OpenStudy (anonymous):

could we try my question?

OpenStudy (anonymous):

2x^3y^4z^3a^2 over 4x^2y^5za Answer choices: A) axz over 2y B) xz^2a over 2y C) ax^2z over 2y D) xy over 2

OpenStudy (anonymous):

which one would it be?

OpenStudy (saranya):

Ans is B) xz^2a over 2y

OpenStudy (kinggeorge):

It should just be \[z^{3-1}=z^2.\]We're doing your question one part at a time. There's a reason I'm choosing these particular variables and exponents. We just have one left to do before we finish. Give a shot at simplifying \[\frac{a^2}{a}\]

OpenStudy (anonymous):

@saranya thanks so much

OpenStudy (anonymous):

@KingGeorge okay thanks

OpenStudy (saranya):

no probs dear :)

OpenStudy (anonymous):

@saranya how did you get this answer it was kinda difficult could you explain please?

OpenStudy (kinggeorge):

She got it using the method I was showing you. We started with a \[\frac{x^3}{x^2}=x\]the last time you asked this, and here, we moved on to get \[\frac{y^4}{y^5}=\frac{1}{y}\qquad \text{and}\qquad \frac{z^3}{z}=z^2\]That means we had just had to simplify \[\frac{a^2}{a}=a\qquad \text{and}\qquad \frac{2}{4}=\frac{1}{2}\]to get the final solution.

OpenStudy (kinggeorge):

You just combine all of those terms to get \[x\cdot\frac{1}{y}\cdot{z^2}\cdot a\cdot\frac{1}{2}=\frac{xz^2a}{2y}\]

OpenStudy (saranya):

yea sure :) It's simple ! Now take the x part 2x^3/4x^2= x/2 Now take the y part y^4/y^5 = 1/y Now take z part z^3/z = z^2 now take a part a^2/a = a Now combine all that is simply multiply all the ans : yu will get xz^2a over 2y

OpenStudy (anonymous):

@KingGeorge thanks so much you helped alot

OpenStudy (anonymous):

thanks to you guys i got an 80 on my test thanks so much

OpenStudy (saranya):

ur welcome dear as always :)

OpenStudy (kinggeorge):

For the future, posting these questions that were on your test is against the code of conduct. Especially if you were only looking for the answer. http://openstudy.com/code-of-conduct